# KnowGPT: Knowledge Injection for Large Language Models

### v4

@@ 4.1.1 Context-relatedness Auxiliary Reward
The key motivation is to encourage paths closely related to the given question context. Specifically, we evaluate the semantic relevance of a path $\mathcal{P}_{i}$ to the context $\mathcal{Q}$. Inspired by the prevailing study Yasunaga et al. (2021), a fixed but well-trained matrix $\boldsymbol{W}$ is applied to map the path embedding $\boldsymbol{\mathcal{P}}$ to the same semantic space with context embedding $\boldsymbol{c}$. To this end, this auxiliary reward is formulated as:
@@
where $\boldsymbol{c}$ is the embedding of context $\mathcal{Q}$ we obtained from a pre-trained LM Kenton and Toutanova (2019) and the embedding of path $\mathcal{P}_{i}$ is the average of the embeddings of all the entities and relations we have walked through till $i$, i.e., $Avg(\boldsymbol{e}_{source}+\boldsymbol{re}_{1}...+\boldsymbol{e}_{i})$, where $i\leq length(\mathcal{P}_{target})$. This step-by-step reward scheme provides rewards before the target is reached.
@@ 4.1.2 Conciseness Auxiliary Reward
There are two additional significant challenges for the candidate reasoning background. $(i)$ The natural limitation of black-box LLMs for over-long context understanding gives constrained budgets for prompts, where the extracted path is expected to be concise enough to ensure the full understanding by black-box LLMs. $(ii)$ The prohibitive cost of calling black-box LLMs’ API guides the prompt to be more concise. By limiting the step size, we encourage the policy to find as much valuable information as possible within the shortest path length.
@@
Finally, we use the trade-off parameters to balance the significance of each reward: $r_{\text{total}}=\lambda_{1}r_{reach}+\lambda_{2}r_{cr}+\lambda_{3}r_{cs}.$, where $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ are hyperparameters.
@@ 4.1.3 Training Policy Network
To solve the MDP defined above, a tailored policy network $\pi_{\theta}(s,a)=p(a|s;\theta)$ is trained to extract a reasoning path in the KG. We optimize the network with policy gradient Xiong et al. (2017). The optimal policy navigates the agent from the source entity to the target entity while maximizing the accumulated rewards.
4.2 Structured Knowledge Prompts
In this subsection, we design a tailored prompt construction strategy based on Multi-Armed Bandit (MAB). The key idea is to learn to select the best path extraction and prompt templates at a meta-level. We will begin by outlining the overall strategy, followed by detailing its instantiation with two path extraction methodologies and three templates.
Suppose we have several path extraction strategies $\{\mathcal{P}_{1},\mathcal{P}_{2},...,\mathcal{P}_{m}\}$ and several candidate formats =$\mathcal{F}=\{\mathcal{F}_{1},\mathcal{F}_{2},...,\mathcal{F}_{n}\}$. Each path extraction strategy $\mathcal{P}_{i}$ is a method for selecting a subgraph given a question context, such as the RL-based strategy discussed above. Every prompt template $\mathcal{F}_{j}$ represents a transformation of the triples within the subgraph into a natural language prompt.
@@ 4.2 Structured Knowledge Prompts
[…]
@@
To capture context-aware correlation between questions and different combinations of knowledge and prompt formats, we formulate the selection mechanism of MAB with an expectation function $E(\cdot)$. It adaptively measures the potential expectation of a combination for different questions.
@@
Empirically maximizing $\boldsymbol{c}\times\bm{\alpha}_{i}$ could encourage exploitation Chen et al. (2019); Dong et al. (2023b) for the best combination, we could effectively update $\bm{\alpha}_{(i)}$ via modeling the correlations between the context embedding of the anchor question ${\bf c}_{i}$ and all the previously selected contexts ${\bf C}_{(i)}$ for particular combination $\mathcal{P}\mathcal{F}_{(i)}$ in former $k$ steps, and the rewards $\textbf{r}_{pf}^{(i)}$ obtained from the selection of the current combination. Concretely, $\bm{\beta}^{(b)}$ is updated as:
@@
Similarly, in order to encourage exploration within less frequently selected pairings, we employ an upper confidence bound approach to balance exploration and exploitation. This is achieved through the introduction of the parameter $\beta^{(i)}$. Inspired by prevailing studies Walsh et al. (2009); Dong et al. (2023b), we can derive the following exploration term $\beta^{(i)}$:
@@ 4.2.1 Implementation
We implement the above MAB strategies with two path extraction strategies and three templates. Note that our MAB design is general and can be implemented with more path extraction strategies and prompt templates for better performance. The path extraction strategies include:
  • [leftmargin=*]
  • $\mathcal{P}_{\text{RL}}$: The RL-based path extraction strategy presented in the previous subsection.
  • $\mathcal{P}_{\text{sub}}$: A heuristic sub-graph extraction strategy that extracts a 2-hop subgraph around both the source and target entities. Detailed implementation can be found in Appendix A.1. Since RL is notoriously unstable Sutton and Barto (2018), we introduce $\mathcal{P}_{\text{sub}}$ as an alternative candidate strategy for the MAB selection, ensuring a fallback option if the RL-based approach does not perform well.
@@
Considering two path extraction methods: $\mathcal{P}_{\text{sub}}$ and $\mathcal{P}_{\text{RL}}$ , as well as three prompt translation methods: $\mathcal{F}_{t}$, $\mathcal{F}_{s}$ and $\mathcal{F}_{g}$, the MAB is trained to learn from the feedback from LLMs to prioritize the most appropriate combination among two extraction methods and three predefined prompt formats for different real-world question contexts, i.e., $\mathcal{P}\mathcal{F}=\{(\mathcal{P}_{sub}\mathcal{F}_{t}),$ $(\mathcal{P}_{sub}\mathcal{F}_{s}),(\mathcal{P}_{sub}\mathcal{F}_{g}),(\mathcal{P}_{RL}\mathcal{F}_{t}),(\mathcal{P}_{RL}\mathcal{F}_{s}),(\mathcal{P}_{RL}\mathcal{F}_{g})\}$.
Table 1: Performance comparison among state-of-the-art baselines and KnowGPT on three benchmark datasets.
@@
*We used ‘text-davinci-002’ provided by OpenAI as the implementation of GPT-3, and ‘gpt-3.5-turbo’ for ChatGPT.
@@
5 Experiments
We conduct extensive experiments to evaluate KnowGPT on three benchmark question-answering datasets, covering both commonsense and domain-specific QA. Our experiments are designed to answer the following research questions:
  • [leftmargin=*]
  • RQ1: How does KnowGPT perform when compared with the state-of-the-art LLMs and KG-enhanced QA baselines?
  • RQ2: Does the proposed MAB-based prompt construction strategy contribute to the performance?
  • RQ3: Can KnowGPT solve complex reasoning tasks, and is KG helpful in this reasoning process?
5.1 Experimental Setup
5.1.1 QA Datasets
We evaluate KnowGPT on three QA datasets spanning two fields: CommonsenseQA Talmor et al. (2019) and OpenBookQA Mihaylov et al. (2018) serve as benchmarks for commonsense reasoning, while MedQA-USMLE Jin et al. (2021) acts as a domain-specific QA benchmark. The statistics of these three datasets can be found in Table 5 in the Appendix.
5.1.2 State-of-the-art Baselines
We carefully select baseline models from four categories for a comprehensive evaluation.
LM + Fine-tuning We compare our method with vanilla fine-tuned LMs. Specifically, we choose Bert-base, Bert-large Kenton and Toutanova (2019), and RoBerta-large Liu et al. (2019) as representative fine-tune LM methods. To conduct commonsense and biomedical QA, we fine-tune these three LMs via an additional linear layer.
KG-enhanced LM We have also implemented several recently released models for integrating KGs into question answering, which encompass MHGRN Feng et al. (2020), QA-GNN Yasunaga et al. (2021), HamQA Dong et al. (2023a), JointLK Sun et al. (2022), GreaseLM Zhang et al. (2022b) and GrapeQA Taunk et al. (2023). To ensure a fair comparison, we implement these baselines with advanced language models that are optimized for particular datasets. Specifically, RoBerta-large Liu et al. (2019) is used for CommenseQA, while AristoRoBERTa Clark et al. (2020) is designated for OpenBookQA. For MedQA, we opt for the top-tier biomedical language model, SapBERT Liu et al. (2021). Note that due to the white-box nature of these methods and their high computation overheads, it is infeasible to apply them to state-of-the-art LLMs, like ChatGPT and GPT-4.
LLM We include several representative generative LLMs, including ChatGLM, ChatGLM2, Baichuan-7B, InternLM, GPT-3, ChatGPT and GPT-4 as knowledge-agnostic alternatives. Specifically, we used the model ‘text-davinci-002’ provided by OpenAI as the implementation of GPT-3, and ‘gpt-3.5-turbo’ and ‘gpt-4’ as the implementations of ChatGPT and GPT-4, respectively (we have provided more implementation details of all LLMs in Appendix A.4). The question-answering task is conducted under the zero-shot setting with the question query from the test set as input.
LLM + KG Prompting To verify the effectiveness of our proposed prompting strategy used KnowGPT, we also add the state-of-the-art KG prompting methods, i.e., CoK Wang et al. (2023), ChatRule Luo et al. (2023), and Mindmap Wen et al. (2023) as baselines.
Table 2: OpenBookQA Official Leaderboard records of three groups of related models (sorted by rankings).
OpenBookQA Leaderboard
Human Performance 0.917
w/o KG 0.778
MHGRN Feng et al. (2020) 0.806
QA-GNN Yasunaga et al. (2021) 0.828
GreaseLM Zhang et al. (2022b) 0.848
HamQA Dong et al. (2023a) 0.850
JointLK Sun et al. (2022) 0.856
GSC Wang et al. (2021) 0.874
UnifiedQA Khashabi et al. (2020) 0.872
DRAGON Yasunaga et al. (2022) 0.878
GenMC Huang et al. (2022) 0.898
GenMC Ensemble Huang et al. (2022) 0.920
MVP-Tuning Ensemble Huang et al. (2023) 0.952
KnowGPT 0.916
@@ 5.2 Main Results (RQ1)
To address RQ1, we evaluate KnowGPT by comparing it to state-of-the-art baselines on the three benchmark datasets. KnowGPT is based on the original ChatGPT. We measure the performance using accuracy, which calculates the percentage of questions correctly predicted by the model out of the total questions. We have following observations:
@@
  • KnowGPT surpasses the performance of ChatGPT and even GPT-4. On average, KnowGPT achieves a 23.7% higher testing accuracy than ChatGPT. Specifically, KnowGPT outperforms ChatGPT by 10.8%, 32.4%, and 29.4% on the CommonsenseQA, OpenBookQA, and MedQA datasets, respectively. More importantly, despite being based on ChatGPT, KnowGPT outperforms the state-of-the-art LLM GPT-4 by 3.3%, 1.4%, and 1.8% on the CommonsenseQA, OpenBookQA, and MedQA datasets, respectively. These results confirm that black-box knowledge injecting can effectively enhance the capabilities of LLMs.
  • KnowGPT outperforms all KG-enhanced LMs significantly. This implies our black-box knowledge injection method proficiently encodes knowledge into LLMs. Furthermore, it showcases the superiority of our model, given its adaptable application to ChatGPT using only the model API, a feat not achievable by white-box methods.
@@ 5.2.1 Leaderboard Ranking
We submit our results onto the official leaderboard maintained by the authors of OpenbookQA. The full records on the leaderboard are shown on the website [^1] https://leaderboard.allenai.org/open˙book˙qa/submissions/public. , while our result can be found from here [^2] https://leaderboard.allenai.org/open˙book˙qa/submission/cj9game4arcuacugbrj0..
We summarize the related submissions in Table 2, including three categories: traditional KG-enhanced LM, fine-tuning of LLMs, e.g., T5-11B used in UnifiedQA, and ensemble of multiple predictors. KnowGPT significantly outperforms traditional KG-enhanced LMs with 4.2% improvements when compared to the best baseline. Despite the 2.6% difference of the state-of-the-art fine-tuning method X-Reasoner, which is the currently best single model, they costly train the T5-11B LLM and also initialize the model on eight A100 GPUs. The third group of methods occupies the leaderboard by leveraging ensemble learning strategies. Nevertheless, KnowGPT can still obtain competitive performance without ensembling with merely 0.4% below GenMC Ensemble Huang et al. (2022). Notably, our KnowGPT is remarkably comparable to the human performance.
Table 3: Ablation study on the effectiveness of two path extraction methods.
Path Extraction Model CSQA OBQA MedQA

### v5

@@ 3.1.1 Context-relatedness Auxiliary Reward
The key motivation is to encourage paths closely related to the given question context. Specifically, we evaluate the semantic relevance of a path $\mathcal{P}_{i}$ to the context $\mathcal{Q}$. Inspired by the prevailing study [69], a fixed but well-trained matrix $\boldsymbol{W}$ is applied to map the path embedding $\boldsymbol{\mathcal{P}}$ to the same semantic space with context embedding $\boldsymbol{c}$. To this end, this auxiliary reward is formulated as:
@@
where $\boldsymbol{c}$ is the embedding of context $\mathcal{Q}$ we obtained from a pre-trained LM [24] and the embedding of path $\mathcal{P}_{i}$ is the average of the embeddings of all the entities and relations we have walked through till $i$, i.e., $Avg(\boldsymbol{e}_{source}+\boldsymbol{re}_{1}...+\boldsymbol{e}_{i})$, where $i\leq length(\mathcal{P}_{target})$. This step-by-step reward scheme provides rewards before the target is reached.
@@ 3.1.2 Conciseness Auxiliary Reward
There are two additional significant challenges for the candidate reasoning background. $(i)$ The natural limitation of LLMs for over-long context understanding gives constrained budgets for prompts, where the extracted knowledge chain is expected to be concise enough to ensure the full understanding by closed-source LLMs. $(ii)$ The prohibitive cost of calling LLMs’ API guides the prompt to be more concise. By limiting the step size, we encourage the policy to find as much valuable information as possible within the shortest path length.
@@
To this end, our overall reward modeling consists of three major criteria that comprehensively incentivize the entire policy learning for an effective knowledge extraction.
@@ 3.1.3 Training Policy Network
To solve the MDP defined above, a tailored policy network $\pi_{\theta}(s,a)=p(a|s;\theta)$ is trained to extract a reasoning chain in the KG. We optimize the network with policy gradient [65]. The optimal policy navigates the agent from the source entity to the target entity while maximizing the accumulated rewards. We provide more training details in the Appendix.
3.2 Prompt Construction with Multi-armed Bandit
In this subsection, we design a tailored prompt construction strategy based on Multi-Armed Bandit (MAB). The key idea is to learn to select the best knowledge extraction and prompt templates at a meta-level. We will begin by outlining the overall strategy, followed by detailing its instantiation with two knowledge extraction methodologies and three templates.
Suppose we have several knowledge extraction strategies $\{\mathcal{P}_{1},\mathcal{P}_{2},...,\mathcal{P}_{m}\}$ and several candidate prompt formats =$\mathcal{F}=\{\mathcal{F}_{1},\mathcal{F}_{2},...,\mathcal{F}_{n}\}$. Each knowledge extraction strategy $\mathcal{P}_{i}$ is a method for selecting reasoning background given a question context, such as the RL-based strategy discussed above. Every prompt template $\mathcal{F}_{j}$ represents a mechanism to transform the triples within the subgraph into a prompt for an LLM prediction.
@@ 3.2 Prompt Construction with Multi-armed Bandit
$\sigma(f(\mathcal{P}\mathcal{F}_{(i)}))=\begin{cases}1&\textit{if }accurate;\\ 0&\textit{otherwise}.\end{cases}$ (5)
@@
To capture the context-aware correlation between questions and different combinations of knowledge and prompt formats, we formulate the selection mechanism of MAB with an expectation function $E(\cdot)$. It adaptively measures the potential expectation of a combination for different questions.
@@
Empirically maximizing $\boldsymbol{c}\times\bm{\alpha}_{i}$ could encourage exploitation [8; 11] for the best combination, we could effectively update $\bm{\alpha}_{(i)}$ via modeling the correlations between the context embedding of the anchor question ${\bf c}_{i}$ and all the previously selected contexts ${\bf C}_{(i)}$ for particular combination $\mathcal{P}\mathcal{F}_{(i)}$ in former $k$ steps, and the rewards $\textbf{r}_{pf}^{(i)}$ obtained from the selection of the current combination. Concretely, the $\bm{\beta}^{(b)}$ is updated as:
@@
Similarly, in order to encourage exploration within less frequently selected pairings, we employ an upper confidence bound approach to balance exploration and exploitation. This is achieved through the introduction of the parameter $\beta^{(i)}$. Inspired by prevailing studies [59; 11], we can derive the following exploration term $\beta^{(i)}$:
@@ 3.2.1 Implementation
We implement the above MAB strategies with two knowledge extraction strategies and three templates. Note that our MAB design is general and can be implemented with more knowledge extraction strategies and prompt templates for better performance. The knowledge extraction strategies include:
  • $\mathcal{P}_{\text{RL}}$: The RL-based knowledge extraction strategy presented in the previous subsection.
  • $\mathcal{P}_{\text{sub}}$: A heuristic sub-graph extraction strategy that extracts a 2-hop subgraph around both the source and target entities. Detailed implementation can be found in Section B.1 of Appendix. Since RL is notoriously unstable [54], we introduce $\mathcal{P}_{\text{sub}}$ as an alternative candidate strategy for the MAB selection, ensuring a fallback option if the RL-based approach does not perform well.
@@
Considering two knowledge extraction methods: $\mathcal{P}_{\text{sub}}$ and $\mathcal{P}_{\text{RL}}$, as well as three prompt translation methods: $\mathcal{F}_{t}$, $\mathcal{F}_{s}$ and $\mathcal{F}_{g}$, the MAB is trained to learn from the feedback from LLMs to prioritize the most appropriate combination among two extraction methods and three predefined prompt formats for different real-world question contexts, i.e., $\mathcal{P}\mathcal{F}=\{(\mathcal{P}_{sub}\mathcal{F}_{t}),$ $(\mathcal{P}_{sub}\mathcal{F}_{s}),(\mathcal{P}_{sub}\mathcal{F}_{g}),(\mathcal{P}_{RL}\mathcal{F}_{t}),(\mathcal{P}_{RL}\mathcal{F}_{s}),(\mathcal{P}_{RL}\mathcal{F}_{g})\}$.
4 Experiments
We conduct extensive experiments to evaluate KnowGPT on three benchmark question-answering datasets, covering both commonsense and domain-specific QA. We implement KnowGPT upon GPT-3.5. Our experiments are designed to answer the following research questions:
  • RQ1 (Main results): How does KnowGPT perform when compared with the state-of-the-art LLMs and KG-enhanced QA baselines?
  • RQ2 (Ablation Study): How does each key component of KnowGPT contribute to the performance?
  • RQ3 (Case study): How could KG help solve complex reasoning tasks? See Appendix C.2.
4.1 Experimental Setup
Datasets. We evaluate KnowGPT on three QA datasets spanning two fields: CommonsenseQA [55] and OpenBookQA [41] serve as benchmarks for commonsense reasoning, while MedQA-USMLE [23] acts as a domain-specific QA benchmark. The statistics of these three datasets can be found in Table 5 in the Appendix.
Baselines. We carefully select baseline models from four categories for a comprehensive evaluation. LM + Fine-tuning. We compare our method with vanilla fine-tuned LMs. Specifically, we choose Bert-base, Bert-large [24], and RoBerta-large [38] as representative fine-tune LM methods. To conduct commonsense and biomedical QA, we fine-tune these three LMs via additional linear layers.
KG-enhanced LM. We have also implemented several recently released models for integrating KGs into question answering, which encompass MHGRN [13], QA-GNN [69], HamQA [10], JointLK [53], GreaseLM [75] and GrapeQA [56]. To ensure a fair comparison, we implement these baselines with advanced language models that are optimized for particular datasets. Specifically, RoBerta-large [38] is used for CommenseQA, while AristoRoBERTa [9] is designated for OpenBookQA. For MedQA, we opt for the top-tier biomedical language model, SapBERT [35]. Note that due to the white-box nature of these methods and their high computation overheads, it is infeasible to apply them to state-of-the-art LLMs, like GPT-3.5 and GPT-4.
LLM + Zero-shot. We include several representative generative LLMs, including ChatGLM, ChatGLM2, Baichuan-7B, InternLM, GPT-3, GPT-3.5 and GPT-4 as knowledge-agnostic alternatives. Specifically, we used the model ‘text-davinci-002’ provided by OpenAI as the implementation of GPT-3, and ‘gpt-3.5-turbo’ and ‘gpt-4’ as the implementations of GPT-3.5 and GPT-4, respectively (we have provided more implementation details of all LLMs in Appendix A.4). The question-answering task is conducted under the zero-shot setting with the question query from the test set as input.
LLM + KG Prompting. To verify the effectiveness of our prompting strategy, we also add the state-of-the-art KG prompting methods, i.e., CoK [60], RoG [39], and Mindmap [63] as baselines. Notably, we did include KGR [18] in our main results, since the authors have not released their codes.
Table 1: Performance comparison among baseline models and KnowGPT on three benchmark datasets.
@@ 4.1 Experimental Setup
*We used ‘text-davinci-002’ and ‘gpt-3.5-turbo’ provided by OpenAI as the implementation of GPT models.
@@ 4.2 Main Results (RQ1)
To address RQ1, we evaluate KnowGPT by comparing it to state-of-the-art baselines on the three benchmark datasets. KnowGPT is based on the original GPT-3.5. We measure the performance using accuracy, which calculates the percentage of questions correctly predicted by the model out of the total questions in the test set. We have the following observations:
@@
  • KnowGPT surpasses the performance of GPT-3.5 and even GPT-4. On average, KnowGPT achieves a 23.7% higher testing accuracy than GPT-3.5. Specifically, KnowGPT outperforms GPT-3.5 by 10.8%, 32.4%, and 29.4% on the CommonsenseQA, OpenBookQA, and MedQA datasets, respectively. More importantly, despite being based on GPT-3.5, KnowGPT outperforms the state-of-the-art LLM GPT-4 by 3.3%, 1.4%, and 1.8% on the CommonsenseQA, OpenBookQA, and MedQA datasets, respectively. These results confirm that black-box knowledge injecting can effectively enhance the capabilities of LLMs.
  • KnowGPT outperforms all KG-enhanced LMs significantly. This implies our black-box knowledge injection method proficiently encodes knowledge into LLMs. Furthermore, it showcases the superiority of our black-box approach, given its adaptable application to GPT-3.5 using only the model API, a feat not achievable by white-box methods.
@@ 4.2.1 Leaderboard Ranking
We submit our results onto the official leaderboard maintained by the authors of OpenbookQA. The full records on the leaderboard are shown on the website [^1] https://leaderboard.allenai.org/open_book_qa/submissions/public. , while our result can be found from here [^2] https://leaderboard.allenai.org/open_book_qa/submission/cp743buq4uo7qe4e9750..
We summarize the related submissions in Table 2, including three categories: traditional KG-enhanced LM, fine-tuning of LLMs, e.g., T5-11B used in UnifiedQA, and ensemble of multiple predictors. KnowGPT significantly outperforms traditional KG-enhanced LMs with 5.2% improvements when compared to the best baseline. The third group of methods occupies the leaderboard by leveraging ensemble learning strategies. Nevertheless, KnowGPT can still obtain competitive performance without ensembling with 0.6% above GenMC Ensemble [22]. Notably, our KnowGPT is remarkably comparable to the human performance.
Table 2: OpenBookQA Official Leaderboard records of three groups of related models.
      Human Performance (0.917)
      w/o KG       0.778       UnifiedQA [25]       0.872
      MHGRN [13]       0.806       DRAGON [68]       0.878
      QA-GNN [69]       0.828       GenMC [22]       0.898
      GreaseLM [75]       0.848       Human Performance       0.917
      HamQA [10]       0.850       GenMC Ensemble [22]       0.920
      JointLK [53]       0.856       X-Reasoner [21]       0.952
      GSC [61]       0.874       KnowGPT       0.926
4.3 Ablation Studies (RQ2)
To answer RQ2, we conduct two ablation studies. First, in Table 4, we measure the importance of the tailored reinforcement learning-based knowledge extraction module, i.e., $\mathcal{P}_{\text{RL}}$. Specifically, we compare it with the heuristic sub-graph extraction strategy, i.e., $\mathcal{P}_{\text{sub}}$. The performance is evaluated by directly feeding the extracted knowledge with the prompt format of ‘Sentence’, i.e., $\mathcal{F}_{s}$, to GPT-3.5. We also include ‘w/o KG’ as the baseline where GPT-3.5 is asked to independently answer the given question with no reasoning background provided. The results clearly indicate the vital role of our proposed knowledge extraction strategies. Second, we compare each of the three prompt formats subject to the same extracted knowledge. The detailed results are shown in Table 4. Though different formats perform similarly within the difference of 2.2% - 3.3%, they are particularly suitable for different kinds of questions. We illustrate this observation in the following case study section. Both ablation studies support the indispensability of each module, armed with a tailored deep reinforcement learning-based knowledge extraction and a context-aware prompt translation, our KnowGPT performs best on all three benchmark datasets.
Table 3: Ablation study on the effectiveness of two knowledge extraction methods.
Knowledge Extraction Model CSQA OBQA MedQA